36. lines through (-2,0) with slope -3/4.
38. line through (0,9) that is parallel to the line 3x + 5y =15.
40. line through (-2,4) that is perpendicular to the x-axis.
42. slope intercept form. y + 2 = -6 (x-1)

Answer by checkley75(3666)

36) THE LINE FORMULA IS:
Y=mX+b WHERE (m)=SLOPE & (b)=Y INTERCEPT.
NOW SUBSTITUTE THE X & Y POINTS & SOLVE FOR THE Y INTERCEPT (b).
0=-3/4*-2+b
0=6/4+b
b=-6/4
b=-1.5 FOR THE Y INTERCEPT. THUS THE LINE EQUATION IS:
Y=-3X/4-1.5 OR Y=-.75X-1.5 ANSWER.
-----------------------------------------------------------------
38) 3X+5Y=15
5Y=-3X+15
Y=-3X/5+15/5
Y=-.6X+3
THUS A LINE PARALLEL TO THIS LINE & PASSING THROUGH (0,9) HAS A SLOPE OF -.6. NOW SUBSTITUTE THE X & Y POINTS (0,9) & SOLVE FOR THE Y INTERCEPT.
9=-.6*0+b
b=9 IS THE Y INTERCEPT.
THUS THE EQUATION FOR THIS LINE IS Y=-.6X+9
--------------------------------------------------------------
2 SOLUTIONS. NOW YOU FINISH THE HOMEWORK. :-)